OpenStudy (anonymous):

The force on a crate increases uniformly from 30.0N to 40.0N as the crate is pushed from the 0.50m mark to the 0.70m mark. What is the total amount of work done.

OpenStudy (mathmate):

The crate moved horizontally or vertically?

OpenStudy (anonymous):

horizontally

OpenStudy (mathmate):

The force F(x) as a function of x is 50x+5 N (from 30 at x=0.5 to 40 at x=0.7) So work done, \[W = \int\limits_{0.5}^{0.7}F(x)dx = \int\limits_{0.5}^{0.7}(50x+5)dx =[25x^2+5x]_{0.5}^{0.7}\]

OpenStudy (anonymous):

answer?

OpenStudy (mathmate):

Are you able to evaluate the integral on the right hand side?

OpenStudy (anonymous):

Yes. Not really supposed to be using calculus to do this problem, though.

OpenStudy (mathmate):

In that case, you just take the average force and multiply by the distance.

OpenStudy (anonymous):

Use W=Fdcos(theta)

OpenStudy (mathmate):

Cos(theta)=1, assuming force is acting in the direction of the displacement. F varies with position uniformly (linearly), so the product Fav.D is the same as integral of F(x)dx

OpenStudy (anonymous):

Indeed. Got it now....thanks!!!

OpenStudy (mathmate):

Great! You're welcome! :)